# More than Carbon: Cradle-to-Grave environmental impacts of GenAI training on the Nvidia A100 GPU

### v1

@@ Discussion
Luccioni et al. (2023) propose an estimate of the climate change impacts of training BLOOM [26]. The study encompasses the impact of the GPU cards during both the training and inference phases, in addition to the impact of the broader data center infrastructures surrounding the GPU cards, including both cradle-to-gate stage and operational phase. To ensure equivalent system boundaries, the following comparison focuses solely on the embodied and usage impacts of the GPU cards themselves in the context of AI training. Although their study is based on primary data for key usage variables, Luccioni et al. (2023) had to assume the embodied impact of the A100 GPUs due to data limitations. They use an estimate of 150 kg CO2eq per card, directly taken from [4]. Our estimate of 141 kg CO2eq per card is within the same order of magnitude, suggesting that these assumptions can serve as reasonable estimates. It should be noted that our estimate also includes the impacts of distributing (12%) and disposing of the cards (1,5%), which are not included in the original study. This suggests that 150 kg CO2eq is likely a conservative estimate. With a lifespan of six years and 85% utilization ratio compared to our three-year lifespan and 100% utilization, the embodied impacts per GPU hour diverge: from 0.003 kg CO2eq in the original study to 0.0054 kg CO2eq in this study. However, when applying the same assumptions, we also converge roughly on an impact per GPU hour of 0.003 kg CO2eq. Their global estimate for the embodied impacts of training BLOOM is 3.64 t CO2eq, compared to our estimate of 5.8 t CO2eq.
Regarding power consumption, Luccioni et al. (2023) also adopt the TDP (400 W) as a proxy for the average power consumption of an A100 [26]. The climate change impact of the card during its use phase is estimated to be 24.69 t CO2eq, which is lower than our estimate (38.9 t CO2eq). The discrepancy can be attributed to the difference in the carbon intensity assumed for the French electricity mix under consideration (see Fig. 9). The original study reports a value of 57 g CO2eq/kWh, whereas the intensity used for this study is 90 g CO2eq/kWh. The difference can be explained by the fact that the source used by Luccioni et al. employs a production mix that does not account for transmission and distribution losses in the network as well as the import and export of the electricity [41] while the CODDE database employed in this study accounts for these factors.
@@
Figure 9: Comparison of total emission results in tCO2 eq. estimated by Luccioni et al (2023) [26] and by this study (author illustration).
@@
Overall, the estimates between the two studies are consistent in terms of both absolute value and distribution between life cycle stages. This convergence is particularly noteworthy when considering the same usage assumption. Our primary data collection approach validates that carbon-related estimates in previous studies are reasonably accurate, which is consistent with the fact that carbon accounting results exhibit less variation between studies due to standardized methodologies and relatively consistent characterization of elementary flows related to climate change across LCA databases. However, including the results from the precise teardown and elemental analysis reveals substantial variations in impact categories beyond climate change that are not captured by available LCA databases. For instance, adjusting the elemental quantities in the LCI based on our elemental analysis increases the impacts from raw material acquisition, processing and manufacturing by an average of 3.93%, with a minimum increase of 0.02% for freshwater eutrophication and a maximum increase of 32.82% for abiotic resource depletion of minerals and metals. For climate change, the difference is a 1.77% increase.
@@
This demonstrates that while carbon footprint estimates remain relatively stable when using secondary versus primary data, other impact categories - particularly those related to resource depletion and toxicity - show substantial differences, highlighting the importance of primary data collection for comprehensive environmental assessments of specific hardware (e.g. GPUs for AI training).
@@
Beyond Carbon Tunnel Vision: Multi-Criteria Environmental Assessment of AI
@@
Our LCA extends beyond carbon as single indicator by examining 16 environmental impact categories, representing the first comprehensive cradle-to-grave multi-criteria environmental assessment of AI training. This approach aims to break out of the ‘carbon tunnel vision’ narrative that dominates headlines when discussing AI sustainability, revealing environmental burdens that remains invisible in carbon-only assessments. However, not all 16 impact categories can be interpreted equally. Most impact categories in LCA treat environmental burdens as spatially uniform or at a very low level of spatial precision, with the notable exception of water use, which includes location-specific characterization factors during the impact assessment phase. This lack of spatial differentiations between local and global impacts creates interpretation challenges when comparing results across categories and obscures targeted action to address environmental concerns. More fundamentally, the impact category climate change is a global concern with well-understood mechanisms, while others - like human toxicity or ecotoxicity - manifest as highly localized impacts and are less intuitive for most stakeholders to interpret, and require more environmental research to properly understand the mechanisms behind the environmental impacts pathways.
@@
For example, training GPT-4 demonstrates striking disparities across life cycle stages and impact categories that illustrate this interpretation challenge. GPU hardware manufacturing processes account for 94.5% of human toxicity, cancer impacts, while representing less than 3% of total climate change impacts. In contrast, the operational phase contributes almost 97% to the climate change category, and only 5.5% to human toxicity, cancer impacts (see Fig. 5). The abstract unit ‘Comparative Toxic Units for humans’ (CTUh) used to measure the impact category human toxicity, cancer, does not convey the scale or distribution of health outcomes, which complicates interpretation efforts. For all non-LCA practitioners; the indicator CTUh expresses the potential carcinogenic risk increase in humans from exposure to carcinogenic substances, representing the estimated increase in morbidity cases in the total human population per unit mass of a substance emitted. However, this cryptic unit makes it difficult to quantify the scale of the underlying environmental impacts, limiting stakeholder comprehension and policy relevance. More critically, since LCA results are not site-specific, it remains nearly impossible to interpret how these cancer impacts actually affect specific populations or determine which exact populations bear these health risks. While manufacturing-related carcinogens likely affect nearby communities, and/or communities near extraction sites, the CTUh units does not indicate whether this affects one person severely or one million people minimally. Recent research has begun to address this spatial blindness by identifying where life cycle stages take place, also called ‘regionalized LCA’.
@@
Component manufacturing-related health impacts are geographically concentrated near semiconductor facilities, which are primarily located in East and Southeast Asia [43, 51], where environmental health protections may be less stringent than in region where AI training occurs. Furthermore, semiconductor fabrication requires significant amounts of water while being concentrated in water-stressed regions experiencing regular droughts [43, 12]. Yet, the populations who suffer from AI’s environmental impacts are often not the same as those benefiting from AI services [11], highlighting inequalities throughout the value chain that remain invisible in conventional environmental assessments.
When environmental justice implications are taken into consideration, the environmental impact interpretation assumes an additional dimension. An environmental justice lens highlights the linkages between environmental and social conditions and sheds light on how different levels of environmental quality and protection contribute to the health and well-being of some groups, while harming the welfare of others [35]. In the context of AI development, environmental justice would then encompass the equitable distribution of environmental benefits and burdens, ensuring that no population bears a disproportionate share of negative environmental consequences from AI hardware production and use. However, the supply chain of AI development reveals inequitable distribution patterns that disadvantage certain populations while privileging others [55, 11]. This extends beyond the category of human toxicity to multiple impact categories that lack spatial distribution differentiation. The environmental impact profile of AI gives rise to a concerning pattern of environmental injustice. Populations in Western countries, which primarily benefit from AI services (or at least have access to them), experience mostly the global, diffuse impacts of the use phase, primarily in the form of climate change effects distributed across the entire planet. In addition, some localized impacts, such as particulate matter, are also experienced, depending on the composition of the country’s electricity mix and the geographical distribution of its population in relation to electricity infrastructures. In contrast, populations in extraction, raw material processing, manufacturing and disposal regions receive minimal benefits from AI services but bear the full burden of both localized environmental degradation from production and disposal phases and global climate impacts from the use phase. This spatial imbalance is further compounded by regulatory disparities. AI related environmental impacts are predominantly concentrated in geographical areas where environmental regulations are less restrictive and enforcement mechanisms are inadequate, creating disproportionate risks for vulnerable populations [35, 29].
This geographic disconnect between impact and benefit illustrate how LCA methodologies are limited to capture global inequalities embedded in the AI supply chain, because elementary flows are typically treated as global rather than spatially differentiated. Additionally, the interpretation of non-climate impact categories faces challenges by mixing midpoint and endpoint assessments creating methodological inconsistencies. Placing human health impacts (aggregated endpoint categories) on the same analytical level as e.g. water consumption (midpoint) lacks strict methodological rigor. Nonetheless, the solution is not to neglect non-carbon impacts, but to develop tools to better account for them. Future research could explore social LCAs that could challenge the common narrative of ‘AI for Good’ or the use of ‘AI for Sustainability’ [10] by revealing how AI development creates environmental winners and losers.
@@
This discussion shows that carbon-only accounting fundamentally misrepresents the globally distributed impact profile associated with AI development, creating a dangerous blind spot in the sustainability discourse. Moving beyond carbon tunnel vision requires not only expanding the scope of environmental indicators but also developing methodologies that can meaningfully address environmental justice concerns in the global AI supply chain. Comprehensive environmental assessment frameworks must incorporate multi-criteria optimization approaches that account for the geographic, temporal, and impact-specific distributions of AI environmental burdens. Ideally, the geographic distribution of toxicity assessments and ecosystem impact evaluation should become a fundamental perspective of future sustainable AI development strategies, although it strongly depends on upcoming necessary improvements in LCA databases and methodologies.
@@
Regarding the usage impact the model is highly sensitive to GPU lifespan, their power consumption and utilization ratio. However, obtaining precise values for these parameters is constrained by both data limitations and their variability across different usage contexts. In other words, they are uncertain, making them important parameters [42] to further investigate. Consequently, it is advisable to consider bounds for the lifespan, maximum power draw and utilization ratio (see SM). In the context of empirical research conducted on an equivalent perimeter with primary usage data, it is imperative to prioritize the collection of these data to ensure optimal quality.
@@
However, a comparison of the cradle-to-gate impact estimates obtained using another background database (imec net zero, [2]) exclusively for the GPU chip (main die and VRAM) highlights significantly different impact results (around 50%) for the climate change criteria. This difference can be explained by modeling assumptions and by differences in the data values provided by the two data set providers, imec net zero and Negaoctet. Therefore, the results cannot be considered as ground truth, but rather depend on the background databases used [37].
@@ Conclusion
The Supplementary Materials will be provided shortly, with anticipated availability in September 2025.

### v2

@@ Results
Luccioni et al. (2023) estimated the climate change impacts of training BLOOM encompassing the impact of the GPU cards during both the training and inference phases, assuming an embodied footprint of 150 kg CO2eq per A100 GPU based on secondary data [luccioni2023]. To ensure equivalent system boundaries, the following comparison focuses solely on the embodied and usage impacts of the GPU cards themselves in the context of AI training. Our cradle-to-grave analysis yields a comparable value of 141 kg CO2eq per GPU. It should be noted that our estimate includes the impacts of distribution (12%) and disposal (1,5%), which are not included by Luccioni et al. (2023), suggesting the previous estimate is conservative. With a lifespan of six years and 85% utilization ratio compared to our three-year lifespan and 100% utilization, the embodied impacts per GPU hour diverge: from 0.003 kg CO2eq in the original study to 0.0054 kg CO2eq in this study. However, under consistent lifetime and utilization assumptions, both studies converge on an embodied impact of approximately 0.003 kg CO2eq per GPU hour. Their global estimate for the embodied impacts of training BLOOM is 3.64 t CO2eq, compared to our estimate of 5.8 t CO2eq.
For the use phase, Luccioni et al. report 24.7 t CO2eq compared to 38.9 t CO2eq in this study, primarily due to differences in the electricity carbon intensity applied (57 g CO2/kWh vs. 90 g CO2/kWh) (see Fig. 9). The difference likely stem from Luccioni et al. employing a production mix that does not account for transmission and distribution losses in the network as well as the import and export of the electricity [FRcarbonIntensity_source] while the CODDE database employed in this study accounts for these factors.
@@
Figure 9: Comparison of total emission results in tCO2 eq. estimated by Luccioni et al (2023) [luccioni2023] and by this study (author illustration).
@@
Despite methodological variations, results align closely in both absolute value and distribution across life cycle stages, validating the robustness of prior carbon-based assessments. Hence, our primary data collection approach validates that carbon-related estimates in previous studies are reasonably accurate. This is consistent with the fact that carbon accounting results exhibit less variation between studies due to standardized methodologies and relatively consistent characterization of elementary flows related to climate change across LCA databases. However, integrating primary teardown and elemental data reveals that non-carbon impact categories differ more substantially. Adjusting the life cycle inventory based on our compositional analysis increased total impacts by an average of 3.9%, with up to 32.8% for abiotic resource depletion of minerals and metals.
@@
This underscores that while carbon estimates remain relatively stable across studies (using secondary versus primary data), other environmental impact categories are more sensitive to primary material data, emphasizing the need for primary data collection and comprehensive, multi-criteria assessments of AI hardware.
@@
Discussion: Beyond Carbon Tunnel Vision - Multi-Criteria Environmental Assessment of AI
@@
This study addresses a critical gap in current AI sustainability discourse by examining environmental impacts beyond carbon emissions that currently dominate both academic discourse and public debate. Our cradle-to-grave LCA of AI model training examines 16 environmental impact categories, representing the first comprehensive multi-criteria environmental evaluation of large-scale AI systems. By systematically quantifying impacts ranging from resource depletion and ecotoxicity to human health effects, we demonstrate that ‘carbon tunnel vision’ fails to capture the full spectrum of environmental consequences associated with AI development. However, not all 16 impact categories can be interpreted equally. Most impact categories in LCA treat environmental burdens as spatially uniform or at a very low level of spatial precision, with the notable exception of water use, which includes location-specific characterization factors during the impact assessment phase. This limited spatial differentiations between local and global impacts complicates direct comparisons across categories and obscures targeted action to address environmental concerns. More fundamentally, the impact category climate change is a global concern with well-understood mechanisms, while others (like human toxicity or ecotoxicity) manifest as highly localized impacts and are less intuitive for most stakeholders to interpret, and require more environmental research to properly understand the mechanisms behind the environmental impacts pathways.
@@
The case of GPT-4 training exemplifies these interpretative challenges through striking disparities across life cycle stages and impact categories. GPU hardware manufacturing accounts for 94.5% of human toxicity (cancer) impacts yet contributes less than 3% to total climate change impacts. In contrast, the operational phase contributes almost 97% to the climate change category but only 5.5% to human toxicity (cancer) (see Fig. 5). The abstract unit ‘Comparative Toxic Units for humans’ (CTUh) used to measure the impact category human toxicity (cancer) does not convey the scale or distribution of health outcomes, which complicates interpretation efforts. For all non-LCA practitioners; the indicator CTUh expresses the potential carcinogenic risk increase in humans from exposure to carcinogenic substances, representing the estimated increase in morbidity cases in the total human population per unit mass of a substance emitted. However, this cryptic unit makes it difficult to quantify the scale of the underlying environmental impacts, limiting stakeholder comprehension and policy relevance. More critically, since LCA results are not site-specific, it remains nearly impossible to interpret how these cancer impacts actually affect specific populations or determine which exact populations bear these health risks. While manufacturing-related carcinogens likely affect nearby communities, and/or communities near extraction sites, the CTUh units does not indicate whether this affects one person severely or one million people minimally.
@@
Recent research has begun to address this spatial blindness by identifying where life cycle stages take place, also called ‘regionalized LCA’. For example, component manufacturing-related health impacts are geographically concentrated near semiconductor facilities, which are primarily located in East and Southeast Asia [roussilhe2022, wang2023semicon_water], where environmental health protections may be less stringent than in region where AI training occurs. Water consumption presents a particularly acute example: semiconductor manufacturing is both water-intensive and geographically concentrated in water-stressed regions experiencing recurrent drought conditions, including Taiwan and other key production hubs [roussilhe2022, TaiwanDrought_Feng2023]. These spatial patterns reveal a fundamental asymmetry in the distribution of AI’s environmental costs and benefits. Populations bearing the health and resource burdens of the hardware life cycle are often not the same as those able to access AI services [falk2024attribution], highlighting inequalities throughout the value chain that remain invisible in conventional environmental assessments.
This geographic decoupling of impacts and benefits raises critical environmental justice concerns that remain obscured in spatially aggregated assessments, underscoring the need for attribution frameworks that can trace environmental responsibility across global AI supply chains and inform equitable governance mechanisms. An environmental justice lens highlights the linkages between environmental and social conditions and sheds light on how different levels of environmental quality and protection contribute to the health and well-being of some groups, while harming the welfare of others [EJ_OECD2024]. In the context of AI development, environmental justice would then encompass the equitable distribution of environmental benefits and burdens, ensuring no population bears disproportionate negative consequences from AI hardware production and use while others capture technological and economic gains. However, current research indicates substantial inequities across AI supply chains: populations in Western and technologically advanced economies, the primary AI service beneficiaries, experience predominantly diffuse global impacts, particularly climate change, along with localized effects such as particulate matter depending on electricity infrastructure. Conversely, marginalized populations in extraction, processing, manufacturing, and disposal regions receive minimal service benefits or access to AI systems while bearing both concentrated local environmental degradation (toxic emissions, resource depletion, ecosystem damage) as well as global climate impacts [zhou2022carbon_Inequal., falk2024attribution]. The resulting impact profile suggests systematic environmental injustice embedded within AI development pathways, where global supply chain configurations externalize environmental costs to communities with limited regulatory protection or political voice creating disproportionate risks for vulnerable populations [EJ_OECD2024, mensah2025missing]. The geographic disconnect between impact and benefit illustrate how LCA methodologies are limited to capture global inequalities embedded in the AI supply chain, because elementary flows are typically treated as global rather than spatially differentiated. Additionally, the interpretation of non-climate impact categories faces challenges by mixing midpoint and endpoint assessments creating methodological inconsistencies. Placing human health impacts (aggregated endpoint categories) on the same analytical level as e.g. water consumption (midpoint) lacks strict methodological rigor. Nonetheless, the solution is not to neglect non-carbon impacts, but to develop tools to better account for them. Future research could explore social LCAs that could challenge the common narrative of ‘AI for Good’ or the use of ‘AI for Sustainability’ [falk2024challenging] by revealing how AI development creates environmental winners and losers.
@@
Carbon-only accounting fundamentally misrepresents AI’s globally distributed impact profile, creating dangerous blind spots in the sustainability discourse. Moving beyond carbon tunnel vision requires not only expanding the scope of environmental indicators but also developing methodologies that can meaningfully address environmental justice concerns across global AI supply chains. Comprehensive environmental assessment frameworks must incorporate multi-criteria optimization approaches that account for geographic, temporal, and impact-specific distributions of environmental burdens. Spatially explicit toxicity and ecosystem impact assessments should become foundational to sustainable AI development strategies, contingent on critical improvements in LCA databases and characterization methods.
@@
Regarding the usage impact the model is highly sensitive to GPU lifespan, their power consumption and utilization ratio. However, obtaining precise values for these parameters is constrained by both data limitations and their variability across different usage contexts. In other words, they are uncertain, making them important parameters [Rosenbaum2018] to further investigate. Consequently, it is advisable to consider bounds for the lifespan, maximum power draw and utilization ratio (see SM). In the context of empirical research conducted on an equivalent perimeter with primary usage data, it is imperative to prioritize the collection of these data to ensure optimal quality.
@@
However, a comparison of the cradle-to-gate impact estimates obtained using another background database (imec net zero, [Boakes2023-jm]) exclusively for the GPU chip (main die and VRAM) highlights significantly different impact results (around 50%) for the climate change criteria. This difference can be explained by modeling assumptions and by differences in the data values provided by the two data set providers, imec net zero and Negaoctet. Therefore, the results cannot be considered as ground truth, but rather depend on the background databases used [su12239948].
@@ Conclusion
You can find the SM here: github.com/sophia-falk/more-than-carbon.